# Erasing Without Remembering: Safeguarding Knowledge Forgetting in Large Language Models

### v1

@@ Abstract
In this paper, we explore machine unlearning from a novel dimension, by studying how to safeguard model unlearning in large language models (LLMs). Our goal is to prevent unlearned models from recalling any related memory of the targeted knowledge. We begin by uncovering a surprisingly simple yet overlooked fact: existing methods typically erase only the exact expressions of the targeted knowledge, leaving paraphrased or related information intact. To rigorously measure such oversights, we introduce UGBench, the first benchmark tailored for evaluating the generalisation performance across 13 state-of-the-art methods. UGBench reveals that unlearned models can still recall paraphrased answers and retain target facts in intermediate layers. To address this, we propose PerMU, a perturbation-based method that significantly enhances the generalisation capabilities for safeguarding LLM unlearning. Experiments demonstrate that PerMU delivers up to a 50.13% improvement in unlearning while maintaining a 43.53% boost in robust generalisation. Our code can be found in https://github.com/MaybeLizzy/UGBench.
@@ 1 Introduction
To further substantiate this observation, our first contribution in this paper is to introduce UGBench, the first generalisation evaluation benchmark tailored for LLM-based machine unlearning. UGBench provides an unbiased assessment of existing methods in forgetting in-scope unlearning samples. The benchmark spans three data domains: two widely-used machine unlearning datasets, TOFU [18] and the Harry Potter books [19, 20], as well as a popular model editing dataset, ZsRE [21]. Evaluations are performed on 13 existing methods using two language models of different scales, Phi-1.3B and LLaMA2-7B. Our empirical analyses on UGBench uncover the following unique findings that have been overlooked by existing research:
@@
These insights suggest target knowledge is not truly erased, leading to risky unlearning due to generalisation failure.
Motivated by the discoveries from UGBench, our second contribution is to shift the focus from the conventional performance race to an equally vital yet unexplored facet: safeguarding LLM-based model unlearning. Our goal is to prevent unlearned models from retaining any memory of the unlearned samples within the defined unlearning scope in Fig. 1, accomplished by enhancing the generalisation capabilities of the unlearned models. However, this ambitious goal is not without challenges. One vanilla approach is to label all rephrased versions of unlearning samples and their paraphrased answers for unlearning. Yet, this method is prohibitively labor-intensive and time-consuming, motivating the development of a novel solution to safeguard unlearning.
To this end, this paper strives to take a pilot step toward safer machine unlearning with strengthened generalisation, by proposing an innovative probability-perturbation unlearning (PerMU) method that, paradoxically, leverages adversarial examples as its foundation. In particular, rather than treating adversarial examples as mere threats, the goal of PerMU is to reverse their role to enable generalised unlearning by perturbing the most vulnerable tokens in the unlearning samples, thereby forcing the model to generate incorrect answers as if it had never been trained on them. To identify these vulnerable tokens for perturbation, we propose a novel metric, termed as MSM, that quantifies the model’s sensitivity to specific tokens with theoretical guarantees.
Through the analysis based on MSM, we observe that subject tokens in the unlearning samples exhibit the highest sensitivity. Driven by this discovery, PerMU injects random noise into the embeddings of subject tokens prior to model processing. This disrupts the model’s ability to recall factual information by shifting the top-ranked tokens in the next-token probability distribution from factual or answer-related terms to those reflecting grammatical or contextual patterns. PerMU then subtracts the original distribution from the perturbed probability distribution. Finally, PerMU minimises the distance between the subtracted distribution and the unlearned model’s distribution, ensuring that the original correct answers are assigned lower probabilities. As such, PerMU simultaneously addresses both causes of the generalisation challenge identified in UGBench as follows:
@@
In sum, our contribution is a novel generalisation evaluation benchmark, UGBench, that assesses the generalisation capabilities of existing methods to forget knowledge within the unlearning scope, as well as a perturbation-based machine unlearning method, PerMU, that uniquely prevents models from remembering associated facts. Comparative experiments with 13 state-of-the-art approaches on 7 partitions across 3 data domains demonstrate that PerMU achieves up to a 50.13% improvement in unlearning and a 43.53% enhancement in robust generalisation, all while maintaining high model utility and superior generation quality.
@@ 2 Related Work
Unlearning Evaluation. Several studies have examined LLM unlearning from different perspectives [26]. Specifically, Patil et al. [22] investigate the effectiveness of typical model editing methods in removing information from model weights. Hong et al. [27] use vocabulary projections to analyze concept vectors through parametric knowledge traces. Yao et al. [28] evaluate seven different unlearning methods on longer-context tasks across three source domains. Meanwhile, Shi et al. [29] assess six desirable properties of unlearned models across eight unlearning methods. Jia et al. [20] focus on the influence of second-order optimization on unlearning. Li et al. [30] examine malicious use scenarios in biosecurity, cybersecurity, and chemical security. Qiu et al. [31] evaluate four distinct unlearning methods for removing highly interconnected data. Du et al. [32] investigate the risk of knowledge leakage after unlearning. The survey by Liu et al. [17] highlights often-overlooked aspects of existing LLM unlearning research and introduces the concept of unlearning scope. Unlike these existing works, UGBench is the first benchmark to evaluate the state-of-the-art methods on their generalisation ability. Notably, the evaluation covers thirteen unlearning methods across seven partitions from three domains, offering a more diverse and challenging scenario.
@@ 3.1 Dataset
TOFU The TOFU dataset comprises 200 diverse synthetic author profiles, each featuring 20 question-answer pairs. It encompasses four subsets—Forget Set, Retain Set, Real Authors, and World Facts—and supports three forgetting settings: Forget01, Forget05, and Forget10, corresponding to the removal of 1%, 5%, and 10% of the data, respectively. Additionally, it provides a paraphrased version of the forget dataset, and we directly use the rephrased questions for testing.
@@
3.3 Thirteen Benchmarked Algorithms.
@@ 4 Benchmark Analysis
As shown in Tab.IV, when tested on the TOFU dataset, although DPO+GDR achieves a superior ROUGE score, even surpassing the retain model, it fails to effectively reduce the Probability score. When tested on LLaMA2-7B using HP dataset (Tab.VI), all methods encounter significant challenges in forgetting the rephrased unlearning samples, with the Probability score showing a gap of up to 43.48% compared to the retain model. Notably, all methods exhibit a Forget ROUGE score that remains above 90%, indicating that they forget almost nothing. A similar trend is observed on ZsRE dataset (Tab.VII). Interestingly, the ICL-based unlearning method achieves the best performance on subject-replaced examples, while other approaches consistently maintain exceptionally high ROUGE, probability, and F1 scores. Moreover, existing unlearning methods encounter significant challenges in forgetting one-hop reasoning examples. Empirical results underscore the limited generalisation ability of existing unlearning methods.
@@
TABLE II: Experimental results for the TOFU dataset using LLaMA2-7B for Forget05 subset. PerMU outperforms the baselines with a relative improvement of up to 50.13% (78.67 $\rightarrow$39.23) in Forget Probability and 14.43% (56.27$\rightarrow$64.39) on Forget Truth Ratio. Meanwhile, PerMU maintains the highest model utility, effectively preserving retained knowledge.
@@
PerMU 33.66 39.23 64.39 83.63 88.24 41.89 91.30 52.57 68.21 89.60 49.77 63.94 64.89 1.78
(ii) The unlearned model are still capable of recalling paraphrased answers during inference. Answers can be expressed in various forms, but existing methods [10], typically focus on training the model to forget only a specific type of answer, neglecting other rephrased versions. We investigate the likelihood of an unlearned model generating paraphrased answers. Experiments are conducted on the TOFU Forget01 dataset using LLaMA2-7B, where we report the average probability of generating paraphrased answers for unlearning samples ($P_{u}$), and rephrased unlearning samples ($P_{r}$). As shown in Tab. III, the unlearned model assigns up to 17.48% probability to rephrased answers when tested on unlearning samples. Furthermore, the probabilities of paraphrased answers on rephrased unlearning samples tend to be assigned even higher values. These results suggest that unlearned models continue to recall paraphrased answers, increasing the potential for reproducing ground truth and posing challenges for generalisation during inference.
TABLE III: The average probability of the model generating a rephrased answer on TOFU Forget01 using LLaMA2-7B.
$P_{u}$ $P_{r}$ $\Delta$
GA 9.45 10.84 1.38
DPO 17.48 17.91 0.42
NPO 10.74 11.98 1.24
TV 13.51 14.66 1.15
WHP 11.46 12.60 1.14
ULD 9.89 10.31 0.42
PerMU 9.06 9.34 0.28
Figure 3: Visualization of the MSM values for each token across all layers using Phi-1.3B. Subject words have brighter colors and exhibit higher MSM, indicating the model’s greater sensitivity to them.
TABLE IV: Experimental results on the Rephrased TOFU dataset using LLaMA2-7B. Notably, the baseline methods struggle to generalise to rephrased unlearning samples. In contrast, PerMU outperforms the baselines by up to 43.53% in Probability and 20.81% in Truth Ratio.
@@
PerMU 26.69 14.75 71.72 65.06 3.14 28.96 30.14 68.26 61.21 2.07 32.72 37.99 65.96 64.80 1.83
5 Proposed Approach: PerMU
To improve the generalisation of unlearned models, we propose a simple perturbation-based machine unlearning method, termed PerMU, which achieves unlearning by simulating adversarial unlearning samples. To identify which tokens to perturb, in this section, we first introduce a Model Sensitivity Metric, MSM, followed by a detailed explanation of the unlearning pipeline for PerMU.
5.1 Model Sensitivity Metric
Proposition 5.1.
To identify the model’s most sensitive token, we conducted experiments to compute the MSM value of each token across all datasets using Phi-1.3B. As shown in Tab. V, the mean MSM value of the subject words is up to 2.66 times higher than that of other words. We further visualize this by normalizing the MSM values of all tokens in each sentence to a range between 0 and 1. As depicted in Fig. 3, the subject words exhibit higher intensity values in the middle layers, indicating that the model is more sensitive to them. This observation aligns with the functionality of the mid-layer modules in generative language models, which primarily recalls facts related to the subject words [38, 42], subsequently influencing the output to reflect memorized properties. Consequently, the model demonstrates a higher sensitivity to subject words.
TABLE V: Comparison between the average MSM values of subject words and the other words. Subject words exhibit higher MSM values than other tokens, indicating greater sensitivity.
Dataset TOFU Harry ZsRE
Subject Words 0.000561 0.003219 0.005961
Other Words 0.000219 0.001708 0.003082
Ratio 2.66 1.86 1.93
5.2 Perturbed Distribution Matching
Building on the observation above, we propose PerMU to achieve unlearning. Specifically, let $x=(x_{1},..x_{sub_{1}}..x_{sub_{k}}..,x_{m})$ be an unlearning sample consisting of $m$ tokens. There exist $k$ subject words in the sentence. To prevent the model from recalling the subject-related facts and to treat the unlearning sample as adversarial, we inject random noise into the subject words in the embedding space prior to inputting them into the model, resulting in the perturbed unlearning sample $x^{\prime}=(x_{1},..x^{\prime}_{sub_{1}}..x^{\prime}_{sub_{k}}..,x_{m})$ where $x^{\prime}_{sub_{i}}$ represents the perturbed subject tokens.
We first input the unlearning sample $x$ into the unlearning model $f_{\theta_{u}}$ to generate the clean-run next-token probability distribution $p(y|x)$, where $y=f_{\theta_{u}}(y|x)$ represents the model’s output. Similarly, the corrupted-run next-token probability distribution $p(y|x^{\prime})$ is obtained by inputting the perturbed sample $x^{\prime}$ into $f_{\theta_{u}}$.
TABLE VI: Experimental results on the Rephrased Harry Potter dataset. PerMU demonstrates effective generalisation to rephrased forget data, achieving relative improvements of up to 17.18% (84.39$\rightarrow$69.89) in Rephrased Forget ROUGE and 19.87% (81.73$\rightarrow$65.49) in Rephrased Forget F1 when tested on LLaMA2-7B.
@@ 5.2 Perturbed Distribution Matching
PerMU 65.21 35.75 63.02 59.11 31.24 56.02 63.20 1.22 71.88 56.36 69.54 69.89 56.59 65.49 82.65 1.27
As illustrated in Fig. 4, in the clean-run probability distribution $p(y|x)$, the top-ranked tokens with the highest probabilities are informative and fact-related. In contrast, for the corrupted-run probability distribution $p(y|x^{\prime})$, the model fails to recognize the subjects or recall related facts, resulting in token generation that relies primarily on context or grammar, with top-ranked tokens being fact-unrelated. Consequently, the corrupted-run probability distribution intuitively simulates a natural unlearning effect, allowing the model to behave as if it had never been trained on the given example. To replicate such a natural unlearning environment, we subtract $p(y|x)$ from $p(y|x^{\prime})$, using a tuning coefficient $C$ to control the strength of forgetting:
$p(Y_{t}|y_{<t})=p(y|x^{\prime})-C\cdot p(y|x),$ (11)
Consequently, the probability of the fact-related tokens are significantly reduced, while the fact-unrelated tokens remain highly ranked, preserving the distribution of irrelevant tokens and thus maintaining the model’s utility.
We then achieve unlearning by fine-tuning $f_{\theta_{u}}$ to match the new, subtracted logit probability distribution $p(Y_{t}|y_{<t})$. For autoregressive text generation, this is decomposed into a step-wise KL divergence [43]:
Algorithm 1 PerMU: Perturbation-based Unlearning
0:  Unlearning sample $x=(x_{1},x_{2},\dots,x_{m})$ with $m$ tokens, target model $f_{\theta_{u}}$, tuning coefficient $C$
0:  Fine-tuned unlearning model $f_{\theta_{u}}$
1:  Identify $k$ subject tokens $\{x_{sub_{1}},\dots,x_{sub_{k}}\}$ in $x$.
2:  Introduce random noise to the subject tokens to obtain the perturbed sample: $x^{\prime}=(x_{1},\dots,x^{\prime}_{sub_{1}},\dots,x^{\prime}_{sub_{k}},\dots,x_{m})$.
7:  Update $f_{\theta_{u}}$ using gradient descent to minimise $L$.
7:  Fine-tuned unlearning model $f_{\theta_{u}}$.
In general, the advantage of PerMU lies in its generality and simplicity. Regarding generality, compared to previous training-based methods [10, 37], PerMU adopts a different approach by achieving unlearning through logit probability manipulation. On one hand, the corrupted subject token embeddings in the first layer prevent the model from recalling any facts related to the unlearning sample across subsequent middle layers. On the other hand, subtracting the clean-run probability distribution causes the answer and answer-related tokens to drop significantly in the probability distribution. Consequently, the unlearned model fails to generate rephrased answers during inference. Moreover, PerMU preserves the distribution of irrelevant tokens, minimizing side effects and maintaining the model’s utility [44]. In terms of simplicity, PerMU requires no additional training of a reinforced model [25, 14] or scope classifier [9], making the training process more efficient.

### v2

# Erasing Without Remembering: Implicit Knowledge Forgetting in Large Language Models

@@ Abstract
In this paper, we investigate knowledge forgetting in large language models with a focus on its generalisation—ensuring that models forget not only specific training samples but also related implicit knowledge. To this end, we begin by identifying a broader unlearning scope that includes both target data and logically associated samples, including rephrased, subject-replaced, one-hop reasoned, and relation-reversed data. To rigorously evaluate generalisation, we introduce UGBench—the first comprehensive benchmark specifically designed to assess the unlearning of in-scope implicit knowledge covering 13 state-of-the-art methods across three datasets. UGBench reveals that unlearned models can still recall paraphrased answers and retain target facts in intermediate layers. This motivates us to take a preliminary step toward more generalised implicit knowledge forgetting by proposing PerMU—a novel probability perturbation-based unlearning paradigm. PerMU simulates adversarial unlearning samples to eliminate fact-related tokens from the logit distribution, collectively reducing the probabilities of all answer-associated tokens. Experiments are conducted on a diverse range of datasets, including TOFU, Harry Potter, ZsRE, WMDP, and MUSE, using models ranging from 1.3B to 13B in scale. The results demonstrate that PerMU delivers up to a 50.40% improvement in unlearning vanilla target data while maintaining a 40.73% boost in forgetting implicit knowledge. Our code can be found in https://github.com/MaybeLizzy/UGBench.
@@ 1 Introduction
To further substantiate this observation, our first contribution in this paper is to introduce UGBench, the first comprehensive benchmark specifically designed to evaluate the unlearning ability of existing methods in forgetting neighbouring implicit knowledge. The benchmark spans three data domains: two widely-used machine unlearning datasets, TOFU [18] and the Harry Potter books [19, 20], as well as a popular model editing dataset, ZsRE [21]. Evaluations are performed on 13 existing methods using two language models of different scales, Phi-1.3B and LLaMA2-7B. Our empirical analyses on UGBench uncover the following unique findings that have been overlooked by existing research:
@@
These insights suggest that the target knowledge is not fully erased, leading to generalisation failure.
Motivated by the findings from UGBench, we highlight a critical challenge in LLM unlearning: enhancing generalisation in forgetting implicit knowledge. However, achieving this goal is not without challenges. One vanilla approach is to identify and label all relevant latent knowledge (e.g., rephrased versions and paraphrased answers) for unlearning. Yet, this process is prohibitively labour-intensive and impractical, motivating the development of a novel solution to enhance generalisation.
To this end, our second contribution is to take a pilot step towards generalised implicit knowledge forgetting by introducing PerMU, a novel probability-perturbation unlearning method that leverages adversarial examples as its foundation. In particular, rather than treating adversarial examples as mere threats, PerMU uses them constructively by perturbing the most vulnerable tokens in the unlearning samples, forcing the model to generate incorrect answers as if it had never been trained on them. To identify these vulnerable tokens, we propose a novel metric, termed as MSM, that quantifies the model’s sensitivity to specific tokens with theoretical guarantees.
Building on the analysis using MSM, PerMU injects random noise into the embeddings of the most sensitive tokens, disrupting the model’s ability to recall factual information. As a result, the top-ranked tokens in the next-token probability distribution shift away from fact-related terms and toward tokens driven primarily by grammatical structure or contextual cues. Since the probability distribution reflects a language model’s internal knowledge [22], directly adjusting the logit distribution to suppress fact-related information offers an intuitive and effective unlearning strategy. PerMU implements this by subtracting the original distribution from the perturbed one, then fine-tuning the model to minimize the distance to this residual distribution—ensuring that correct fact-related tokens are assigned lower probabilities, thereby achieving unlearning. As such, PerMU simultaneously addresses both causes of the generalisation challenge identified in UGBench as follows:
@@
In sum, our contributions are twofold: (1) a novel evaluation benchmark, UGBench, designed to assess the generalisation capability of existing methods in forgetting implicit knowledge; and (2) an unlearning method based on probability perturbation, PerMU, that effectively prevents models from recalling associated facts. We evaluate PerMU across five data domains, including the WMDP [23] and MUSE [24] datasets, using models of varying scales (1.3Bsimilar-$\sim$13B). Comparative experiments show that PerMU achieves up to a 50.40% improvement in unlearning and a 40.73% enhancement in generalisation, all while maintaining high model utility and superior generation quality.
@@ 2 Related Work
Unlearning Evaluation. Several studies have examined LLM unlearning from different perspectives [29, 30, 31, 32, 33]. Specifically, Patil et al. [25] investigate the effectiveness of typical model editing methods in removing information from model weights. Hong et al. [34] use vocabulary projections to analyze concept vectors through parametric knowledge traces. Yao et al. [35] evaluate seven different unlearning methods on longer-context tasks across three source domains. Meanwhile, Shi et al. [24] assess six desirable properties of unlearned models across eight unlearning methods. Jia et al. [20] focus on the influence of second-order optimization on unlearning. Li et al. [23] examine malicious use scenarios in biosecurity, cybersecurity, and chemical security. Qiu et al. [36] evaluate four distinct unlearning methods for removing highly interconnected data. Du et al. [37] investigate the risk of knowledge leakage after unlearning. The survey by Liu et al. [17] highlights often-overlooked aspects of existing LLM unlearning research and introduces the concept of unlearning scope. In contrast, UGBench offers a more precise and formalized definition of unlearning scope, marking a fundamental distinction from prior work. While Patil et al. [31] explore extraction attacks to recover “deleted” information from intermediate hidden states, their focus lies primarily on model editing techniques such as ROME [38]. In comparison, UGBench provides a comprehensive evaluation of the generalisation ability of existing machine unlearning methods across multiple dimensions. Regarding the generalisation evaluation, Dang et al. [32] focuses on improving generalised model utility, by introducing small Gaussian noise to retain data. Unlike prior works, we aim to enhance the generalisation in unlearning implicit knowledge associated with the target fact.
@@ 3.1 Dataset
TOFU The TOFU dataset comprises 200 diverse synthetic author profiles, each featuring 20 question-answer pairs. It encompasses four subsets—Forget Set, Retain Set, Real Authors, and World Facts—and supports three forgetting settings: Forget01, Forget05, and Forget10, corresponding to the removal of 1%, 5%, and 10% of the data, respectively. Additionally, it provides a paraphrased version of the forget dataset, and we directly use them for testing.
@@ 3.2 Evaluation Metrics
3.3 Thirteen Benchmarked Algorithms
@@ 4 Benchmark Analysis
As shown in Tab.IV, when tested on the TOFU dataset, although DPO+GDR achieves a superior ROUGE score, even surpassing the retain model, it fails to effectively reduce the Probability score. When tested on LLaMA2-7B using HP dataset (Tab.V), all methods encounter significant challenges in forgetting the rephrased unlearning samples, with the Probability score showing a gap of up to 43.48% compared to the retain model. Notably, all methods exhibit a Forget ROUGE score that remains above 90%, indicating that they forget almost nothing. A similar trend is observed on ZsRE dataset (Tab.VI). Interestingly, the ICL-based unlearning method achieves the best performance on subject-replaced examples, while other approaches consistently maintain exceptionally high ROUGE, probability, and F1 scores. Moreover, existing unlearning methods encounter significant challenges in forgetting one-hop reasoning examples. Empirical results underscore the limited generalisation ability of existing unlearning methods.
@@
(ii) The unlearned model are still capable of recalling paraphrased answers during inference. Answers can be expressed in various forms, but existing methods [10], typically focus on training the model to forget only a specific type of answer, neglecting other rephrased versions. We investigate the likelihood of an unlearned model generating paraphrased answers. Experiments are conducted on the TOFU Forget01 dataset using LLaMA2-7B, where we report the average probability of generating paraphrased answers for unlearning samples ($P_{u}$), and rephrased unlearning samples ($P_{r}$). As shown in Tab. II, the unlearned model assigns up to 17.48% probability to rephrased answers when tested on unlearning samples. Furthermore, the probabilities of paraphrased answers on rephrased unlearning samples tend to be assigned even higher values. These results suggest that unlearned models continue to recall paraphrased answers, increasing the potential for reproducing ground truth and posing challenges for generalisation during inference.
TABLE II: The average probability of the model generating a rephrased answer on TOFU Forget01 using LLaMA2-7B.
$P_{u}$ $P_{r}$ $\Delta$
GA 9.45 10.84 1.38
DPO 17.48 17.91 0.42
NPO 10.74 11.98 1.24
TV 13.51 14.66 1.15
WHP 11.46 12.60 1.14
ULD 9.89 10.31 0.42
PerMU 9.06 9.34 0.28
5 Proposed Approach: PerMU
Driven by the insights from UGBench, to achieve our goal of more generalised unlearning, simply reversing the gradient descent loss on forget samples—a common practice—is insufficient to fully erase the memory of implicit knowledge. Recent studies have shown that LM’s internal knowledge is largely captured in its probability distribution [22], which has motivated state-of-the-art unlearning approaches like ULD [14]. These methods suppress fact-related information by modifying the logit distribution: an assistant model is fine-tuned to overfit the ground truth, producing sharp logits that are then subtracted from the original logits to reduce the probability of correct responses.
However, this fine-tuning process narrowly targets ground truth answers while neglecting paraphrased or semantically related alternatives. Consequently, these related tokens may still remain in the subtracted logits, posing inference-time risks of generating semantically equivalent answers. Moreover, these methods are computationally expensive due to additional training and rely heavily on the quality of the assistant model, limiting their scalability and practicality in real-world applications.
Motivated by these limitations, we propose PerMU, a more robust and efficient logit-adjustment method for achieving generalised unlearning. Unlike prior approaches that rely on assistant models, PerMU generates adversarial unlearning samples to jointly suppress all answer-related tokens in the logit space. This is achieved through computationally efficient perturbations that produce a modified distribution, which is then used for logit subtraction and matching. To effectively identify which tokens to perturb, we introduce a theoretically grounded model sensitivity metric, MSM, and present the complete unlearning pipeline in the following subsection.
5.1 Model Sensitivity Metric
Proposition 5.1.
TABLE III: Experimental results for the TOFU dataset using LLaMA2-7B for Forget05 subset. PerMU outperforms the baselines with a relative improvement of up to 50.40% (78.67 $\rightarrow$39.02) in Forget Probability and 11.3% (56.27$\rightarrow$62.63) on Forget Truth Ratio. Meanwhile, PerMU maintains the highest model utility, effectively preserving retained knowledge.
@@ 5.1 Model Sensitivity Metric
PerMU 34.91 39.02 62.63 82.78 84.67 42.21 91.50 50.27 65.18 88.75 50.22 59.60 63.51 1.72
$\textsc{PerMU}^{\dagger}$ 33.66 39.23 64.39 83.63 88.24 41.89 91.30 52.57 68.21 89.60 49.77 63.94 64.89 1.78
5.2 Perturbed Distribution Matching
Building on MSM, we identify the top-$K$most sensitive tokens and inject random noise into their embeddings before feeding them into the model, resulting in the perturbed unlearning sample $x^{\prime}$. We first pass the original unlearning sample $x$ into the unlearning model $f_{\theta_{u}}$ to obtain the clean-run next-token probability distribution $p(y|x)$, where $y=f_{\theta_{u}}(y|x)$ represents the model’s output. Similarly, the corrupted-run next-token probability distribution $p(y|x^{\prime})$ is obtained by feeding the perturbed sample $x^{\prime}$ into $f_{\theta_{u}}$. As illustrated in Fig. 3, the clean-run probability distribution $p(y|x)$ assigns the highest probabilities to tokens that are informative and fact-related. In contrast, for the corrupted-run distribution $p(y|x^{\prime})$, the model fails to recall related facts, resulting in token generation that relies primarily on context or grammar, with top-ranked tokens being fact-unrelated.
Consequently, the corrupted-run probability distribution intuitively simulates a natural unlearning effect, allowing the model to behave as if it had never been trained on the given example. To replicate such a natural unlearning environment, we subtract $p(y|x)$ from $p(y|x^{\prime})$, using a tuning coefficient $C$ to control the strength of forgetting: $p(Y_{t}|y_{<t})=p(y|x^{\prime})-C\cdot p(y|x).$ This subtraction suppresses the probabilities of fact-related tokens, while maintaining high rankings for fact-unrelated tokens. As a result, the distribution of irrelevant tokens is preserved, thereby maintaining the model’s utility.
We then achieve unlearning by fine-tuning $f_{\theta_{u}}$ to match the subtracted logit probability distribution $p(Y_{t}|y_{<t})$. For autoregressive text generation, this is decomposed into a step-wise KL divergence [47]:
Algorithm 1 PerMU: Perturbation-based Unlearning
0:  Unlearning sample $x$ with $m$ tokens, target model $f_{\theta_{u}}$, tuning coefficient $C$
0:  Fine-tuned unlearning model $f_{\theta_{u}}$
1:  Identify the top-$K$most sensitive tokens in $x$ using MSM.
2:  Introduce random noise to the $K$tokens embeddings to obtain the perturbed sample: $x^{\prime}$.
7:  Update $f_{\theta_{u}}$ using gradient descent to minimise $L$.
7:  Fine-tuned unlearning model $f_{\theta_{u}}$.
In general, the advantage of PerMU lies in its generality and simplicity. Regarding generality, on the one hand, the corrupted token embeddings in the first layer prevent the model from recalling any facts related to the unlearning sample across subsequent middle layers. On the other hand, subtracting the clean-run probability distribution causes both the answer and answer-related tokens to drop significantly in the probability distribution. Consequently, the unlearned model fails to generate rephrased answers during inference. Moreover, PerMU preserves the distribution of irrelevant tokens, minimizing side effects and maintaining the model’s utility [48]. In terms of simplicity, the identification of sensitive tokens is automated and PerMU requires no additional training of a reinforced model [28, 14] or scope classifier [9], making the training process more efficient.
TABLE IV: Experimental results on the Rephrased TOFU dataset using LLaMA2-7B. Notably, the baseline methods struggle to generalise to rephrased unlearning samples. In contrast, PerMU outperforms the baselines by up to 40.73% (53.37 $\rightarrow$ 31.63) in Probability and 15.67% (60.75 $\rightarrow$ 70.27) in Truth Ratio.
@@ 5.2 Perturbed Distribution Matching
PerMU 27.19 17.91 70.27 64.30 2.85 29.63 31.63 63.09 63.51 2.07 33.56 40.91 62.70 64.40 1.73
$\textsc{PerMU}^{\dagger}$ 26.69 14.75 71.72 65.06 3.14 28.96 30.14 68.26 61.21 2.07 32.72 37.99 65.96 64.80 1.83
TABLE V: Experimental results on the Rephrased Harry Potter dataset. PerMU demonstrates effective generalisation to rephrased forget data, achieving relative improvements of up to 20.36% (84.39 $\rightarrow$ 67.20) in Rephrased Forget ROUGE and 22.72% (81.73 $\rightarrow$ 63.16) in Rephrased Forget F1 when tested on LLaMA2-7B.
@@
PerMU 56.52 31.41 59.53 54.25 29.22 54.98 62.10 1.30 70.15 55.17 68.85 67.20 54.30 63.16 82.49 1.31
$\textsc{PerMU}^{\dagger}$ 65.21 35.75 63.02 59.11 31.24 56.02 63.20 1.22 71.88 56.36 69.54 69.89 56.59 65.49 82.65 1.27